# VideoGen-of-Thought: A Collaborative Framework for Multi-Shot Video Generation

### v1

@@ 3.2 Problem Definition
Given an one-sentence user input $S$ containing $N$ number of shots, e.g., A set of one-sentence prompts, 30 shots, describe a story of a classic American woman Mary’s life, from birth to death, our method aims to generate a multi-shot, reasonable, and consistent video $V$ with minimal manual intervention. The key challenges are defined as:
  • Multi-Shot: Generating minute-level video containing plenty of shots.
  • Reasonability: Ensuring the generated video maintains a logical narrative and storyline.
  • Consistency: Maintaining temporal and identity consistency across different shots.
Our proposed solution, VideoGen-of-Thought (VGoT), starts from preparing the script $\left\{p_{i}\right\}^{N-1}_{i=0}$ for each shot, drafting keyframes $\left\{I_{i}\right\}^{N-1}_{i=0}$ with identity-preserving, then generating video latents $\{z^{f}\}_{f_{i}=0}^{F-1}$ with a cross-shot smooth mechanism within $N$ shots. VGoT provides a novel approach to overcoming the limitations of existing multiple shots video generation methods.
@@ 4 Method: VideoGen-of-Thought
In this section, we introduce each module in our proposed framework, VideoGen-of-Thought (VGoT), including Script Generation Module, Keyframe Generation Module, Shot-Level Video Generation Module, and Cross-Shot Smooth Module. By breaking down the complex task of long video generation into a series of smaller, well-defined modules, VGoT systematically addresses the key challenges of video length, logic, and cross-shot consistency. This modular decomposition allows each component to focus on specific aspects of the problem, resulting in long, reasonable, and consistent video results.
4.1 Module 1: Script Generation
The script generation task is modeled as a sequence generation problem using a pretrained large language model (LLM) [1, 15, 4]. The process starts with a short one-sentence user prompt $S$ with the number of shots $N$ defined, e.g., a set of one-sentence prompts, 30 shots, describe a story about three avatars (two friends Mike and Jane, one bad guy Tom) and their Fantasy Adventure in an ancient pyramid. We leverage an LLM to prepare short shot descriptions $S^{\prime}=\left\{s_{i}\right\}_{i=0}^{N-1}$ for each shot, which builds the foundation logic for the whole story. The script $p_{i}$ for one shot can be fulfilled from $s_{i}$ as:
$p_{i}(p_{cha},p_{b},p_{r},p_{cam},p_{h})=\mathcal{M}_{\text{LLM}}(s_{i},i)$ (4)
where $p_{cha}$ represents the character description, $p_{b}$ is the background description, $p_{r}$ describes the relation between characters and elements in the scene, $p_{cam}$ describes the camera pose, and $p_{h}$ defines the HDR lighting details. An example is shown in the left top of Figure 3. The LLM is conditioned on both the current $s_{i}$ and the previously generated prompts $p_{i-1}$ to maintain narrative coherence and avoid abrupt transitions between scenes. The algorithm for script generation is iterative, producing a prompt for each video shot as shown in Algorithm 1.
Algorithm 1 Script Generation for Long Video
1:User input $S$, total shot number $N$
2:Initialize $P\leftarrow\emptyset$, $p_{tmp}\leftarrow\emptyset$
3:Initialize Story $S^{\prime}=\{s_{1},s_{2},\dots,s_{N}\}$ with $S$
4:for each short shot description $s_{i}\in S$ do
5:     $p_{i}\leftarrow\mathcal{M}_{\text{LLM}}(s_{i},i)$
6:     Append $p_{i}$ to $P$
7:     $p_{tmp}\leftarrow p_{i}$
8:end for
9:return Prompt Sequence $P=\{p_{1},p_{2},\dots,p_{N}\}$
4.2 Module 2: Keyframe Generation
The keyframes generation task is based on the scripts prepared in the Script Module. For each shot, we use a pre-trained CLIP [47] encoder $E_{\text{text}}$ to obtain text embedding $e^{T}_{i}=E_{\text{text}}(p_{i})$. We leverage the knowledge of LLM used in script module to create some avatar descriptions for this story based on short shot descriptions $S^{\prime}$. For instance, given the $S^{\prime}$ of Mary’s life provided in Sec 3.2, LLM provides the avatar prompts $P_{a}=[P_{a1},P_{a2},\dots]$ of Child Mary, Teenager Mary, Mid-aged Mary, Elder-aged Mary, and Old Mary from five domains defined in Eq 4, and specify each shot prompt with a corresponding avatar.
As introduced in section 3, we use pretrained text-to-image diffusion models $\mathcal{M}_{I}$ [49, 59] to generate $I_{a}=[\mathcal{M}_{I}(P_{a1}),\mathcal{M}_{I}(P_{a2}),\dots]$ representing the avatar appearance from $P_{a}$ through cross-attention [60]. Then we employ IP-Adapter [72] to decouple the key $K$ and value $V$ in the cross-attention in $\mathcal{M}_{I}$ by introducing an additional block that involves the identity-preserving embedding from $I_{a}$. This mechanism ensures that each generated keyframe not only matches the given prompt but also maintains visual consistency with the given avatar through copied $K^{\prime}$ and $V^{\prime}$. The generation of each keyframe can be formalized as:
$I_{i}=\mathcal{M}_{I}(e^{T}_{i},e^{I}_{j})$ (5)
where $e^{T}_{i}$ is the text embedding from $p_{i}$ as mentioned above, $e^{I}_{j}$ is the image embedding of the $j$th IP avatar selected for $i$th shot which is obtained by a CLIP vision model [47]. This approach ensures keyframes $\left\{I_{i}\right\}^{N-1}_{i=0}$ maintain visual consistency in the visual appearance of characters and backgrounds style for the same avatar in different shots.
Figure 3: Visual comparison of VGoT and baselines
@@ 4.2 Module 2: Keyframe Generation
4.3 Module 3: Shot-level Video Generation
Through Script Module and KeyFrame Module, we have already got the scripts and keyframes required by each shot-level video. We employ DynamiCrafter [68] as a conditional video model $\mathcal{M}_{V}$ that is capable of taking both the keyframes and the corresponding textual prompts as input to generate a series of latents $Z_{i}=[z_{1},z_{2},...,z_{k}]$ representing the $i$th shot video clip $V_{i}$ which includes $k$ frames, and this process can be formulated as:
$Z_{i}=\mathcal{M}_{\text{V}}(e^{T}_{i},e^{I}_{i},\epsilon),$ (6)
where the $e^{T}_{i}$ and $e^{I}_{i}$ are the text embedding and image embedding of shot $i$, $Z_{i}\in\mathbb{R}^{f\times h\times w\times d}$ is a series of latent code that can be decoded into the video sequence $V_{i}$ containing $f$ frames generated for shot $i$, and $\epsilon_{i}$ is a noise map with same shape to initialize $Z_{i}$ Note that we use short shot description $s_{i}$ but not the detailed script $p_{i}$ to obtain the $e^{T}_{i}$ , because we found in our experiments that shot-level videos $V_{i}$ will be less dynamic when given prompts that are too complex even though we obtain much better keyframe through $p_{i}$. This phenomenon may have been caused by our video model $\mathcal{M}_{\text{V}}$ receiving simpler textual descriptions during training. As future work, this may be improved if $\mathcal{M}_{\text{V}}$ is trained with detailed scripts like Eq 4.
By incorporating both the keyframe and the textual description to obtain latents $z$ in this video generation process, we secure the generation of high-quality, smooth video clips in the subsequent module.
Table 1: Quantitative comparison with state-of-the-art T2V baselines. We compare average CLIP scores from five domains defined in Eq 4, and the average FC and SC scores within and across shots between VGoT and baseline models. We use $\mathbb{bold}$ to highlight the highest and $\underline{underline}$ for the second high.
Model CLIP (pcha) $\uparrow$ CLIP (pb) $\uparrow$ CLIP (pr) $\uparrow$ CLIP (pcam) $\uparrow$ CLIP (ph) $\uparrow$ FC (Within-Shot) $\uparrow$ FC (Cross-Shot) $\uparrow$ SC (Within-Shot) $\uparrow$ SC (Cross-Shot) $\uparrow$
EasyAnimate [69] 0.4086 0.2429 0.1633 0.1130 0.0722 0.4705 0.0268 0.7969 0.2037
CogVideo [71] 0.4113 0.2432 0.1632 0.1122 0.0701 0.6099 0.0222 0.7424 0.2069
VideoCrafter1 [8] 0.4365 0.2417 0.1535 0.1032 0.0651 0.3706 0.0350 0.7623 0.1867
VideoCrafter2 [9] 0.4000 0.2511 0.1654 0.1140 0.0694 0.5569 0.0686 0.7981 0.1798
VGoT 0.4086 0.2429 0.1633 0.1130 0.0722 0.8138 0.2688 0.9717 0.4276
@@ 4.3 Module 3: Shot-level Video Generation
4.4 Module 4: Smooth Mechanism
The smoothing mechanism ensures a seamless transition from the current shot $i$ to the next shot $i+1$, ultimately achieving overall coherence throughout the final video. Inspired by FIFO [33] that uses latent partitioning to divide latents $Z$ into several parts to support different noise levels across frames, we implement a simple but efficient strategy to secure cross-shot video generation. Since adjacent shots could be dissimilar, we consider it necessary to reset the noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ for $\{z^{f}\}_{f=k\times i}^{(k\times(i+1)}$ generated with Eq 6 at shot $i$, and the length of the reset boundary should be equal to the frame number $k$ for each shot. Each reset boundary contains the conditional embedding $e^{T}_{i}$ and $e^{I}_{i}$, and a series of noise $\epsilon$ set with time step scheduler $\beta_{i}$ as defined in Sec 3. Until the final unit is finished, we can decode the whole $\{z^{f}\}_{f=0}^{F-1}$ into a final video $V$ with $N$ shots, and a visual demonstration of our generated results is shown in Fig 4.
@@ 4.4 Module 4: Smooth Mechanism
CLIP average $\uparrow$ PSNR $\uparrow$ IS $\uparrow$ FC (Within-Shot) $\uparrow$ FC (Cross-Shot) $\uparrow$ SC (Within-Shot) $\uparrow$ SC (Cross-Shot) $\uparrow$
w/o EP w/o IP 0.1146 24.3265 7.4624 0.7364 0.1129 0.9406 0.3650
w EP w/o IP 0.1146 24.3265 7.5783 0.7305 0.1174 0.9471 0.3663
w/o EP w IP 0.1223 23.9228 7.4521 0.8745 0.3291 0.9486 0.4186
Full Model 0.1111 25.7857 7.5194 0.8303 0.2738 0.9487 0.3859
@@ 5.1 Experiment Settings
Datasets
Since current datasets can’t satisfy our requirements for cross-shot story, as defined in Sec 3.2, we use VGoT to create 10 stories, each containing 30 shots, with a user input $S$ and 30 short shot scripts $S^{\prime}$, and a detailed 30 shot scripts $P$ from five domains as Eq 4, making a total of 300 video shots for comparison.
Metrics
In this work, we mainly utilize three metrics for quantitative evaluation: CLIP score [47], PSNR [17], IS [5], Face Consistency (FC) score and Style Consistency (SC) score. FC score is calculated by similarities between face features captured by InsightFace [48, 24, 21, 3, 2, 13, 14, 23, 11, 12] in generated videos, and we obtain SC score by style feature similarities from VGG19 [52]. Specifically, we calculate FC score and SC score from both within one shot and cross shots to evaluate the face identity and visual style consistency.
Implementation Details
Our experiments were conducted on A800 machines. For module implementation, we employed GPT-4o [1] as LLM, Kolor [59] as text-to-image diffusion model $\mathcal{M}_{I}$, DynamiCrafter [68] as conditional video model $\mathcal{M}_{V}$. For comparison, we also implement EasyAnimate [69], CogVideo [30], VideoCrafter1 [8],VideoCrafter2 [9].
5.2 Quantitative Evaluation
In this section, we use the 10 stories defined in Sec 5.1 as dataset, and provide the one-sentence $S$ for VGoT and the detailed scripts of 30 shots for the baselines as input. Then we use the 300 shots for each method to calculate CLIP score from five domains, together with FC score and SC score within and across shots. The quantitative comparison is shown in Table 1. Our method significantly outperforms the baselines in terms of both FC and SC, particularly when evaluated across shots, demonstrating that VGoT effectively maintains character identity and style consistency throughout the entire video, providing seamless transitions. It is also worth noting that the CLIP Scores for our method are not obviously comparable with the baselines. This phenomena is due to the diversity of the generated results, as evidenced by our ablation studies in Sec 5.5, which indicate that VGoT tends to produce richer, more varied content that may not strictly match the prompt but adds to the overall narrative depth and coherence.
5.3 Qualitative Evaluation
In this section, we use the one-sentence user input $S$ for VGoT, and detailed prompts $P$ generated by Script Module as input for baselines, to compare the visual effect of generated video results. Here we take the story “A set of one-sentence prompts, 30 shots, describe a story of a classic American woman Mary‘s life, from birth to death.” as an example, the comparison of results is demonstrated in Fig 3. Our method obviously achieves better visual consistency from style to identity over baselines.
5.4 Human Evaluation
In this section, we present a comprehensive user study comparing the generated results from VGoT and four baseline models. We use 10 stories, each consisting of 30 shots, as input and provide 10 users with a total of 50 multi-shot videos. Each user is shown 10 multi-shot videos selected randomly from the generated content and asked to evaluate them based on three criteria: within-shot consistency, cross-shot consistency, and visual quality, using a rating scale of good, normal, or bad.
The results, presented in Table 2, demonstrate that the videos produced by VGoT are preferred by users, particularly excelling in cross-shot consistency, highlighting the superiority of our method as acknowledged through human evaluation.

### v2

@@ 3.2 Problem Definition
Given a one-sentence user input $S$ specifying $N$ shots (e.g., ”A story of Mary’s life from birth to death”), we aim to generate a multi-shot video $V$ with reasonable storylines and visual consistency from one sentence through minimal manual intervention. The core challenges are:
  • Reasonability: Maintaining logical narrative flow across evolving storylines
  • Consistency: Preserving high-level identity [55, 57, 62] while allowing trait variations (e.g., aging expressions, contextual relationships) across shots
  • Multi-Shot Generation: Producing minute-level videos with diverse yet interconnected shots
VideoGen-of-Thought (VGoT) addresses these through script preparation $\{p_{i}\}^{N-1}_{i=0}$, identity-preserved keyframes $\{I_{i}\}^{N-1}_{i=0}$, and cross-shot video latents $\{z^{f}\}^{F-1}_{f_{i}=0}$. Our framework overcomes limitations in existing multi-shot methods through narrative-visual coherency mechanisms.
@@ 4 Method: VideoGen-of-Thought
Our methodology introduces a structured, step-by-step framework for generating multi-shot videos with reasonable storylines and visual consistency from one sentence with minimal manual intervention, addressing the core challenges of narrative fragmentation, visual inconsistency, and transition artifacts through four distinct yet collaborative modules (Fig 2):
4.1 Dynamic Storyline Modeling with Self-Vaildation
We formulate narrative generation as constrained multi-shot decomposition with auto-regressive validation. Given user prompt $S$ and shot count $N$, our system first generates a story draft $S^{\prime}=\{s_{i}\}_{i=1}^{N}$ of short shot descriptions, then produces structured scripts through:
$\mathcal{P}=\{p_{i}\}_{i=1}^{N}=\bigcup_{i=1}^{N}\mathcal{M}_{\text{LLM}}(s_{i}|\mathcal{C}_{\text{film}},\{p_{j}\}_{j=1}^{i-1})$ (6)
where $\mathcal{C}_{\text{film}}\triangleq(p_{\text{cha}},p_{b},p_{r},p_{\text{cam}},p_{h})$ encodes five cinematic dimensions: $p_{\text{cha}}$ governs character appearance evolution and role relationships, $p_{b}$ ensures background consistency across scene transitions, $p_{r}$ maintains interaction patterns and event causality, $p_{\text{cam}}$ specifies shot composition through camera movements, and $p_{h}$ regulates HDR lighting continuity. The self-validation mechanism employs two criteria:
$\mathcal{V}(p_{i})=\mathbb{I}[C(p_{i},p_{i-1})>\tau_{c}]\cdot\mathbb{I}[K(p_{i},\mathcal{C}_{\text{film}})>\tau_{k}]$ (7)
where $C:\mathcal{P}\times\mathcal{P}\rightarrow[0,1]$ measures narrative coherence through pretrained textual feature extractor $E_{\text{GLM}}$ [18] to compute semantic similarity between consecutive shots, and $K:\mathcal{P}\times\mathcal{C}\rightarrow\{0,1\}$ verifies constraint completeness via rule-based checks against $\mathcal{C}_{\text{film}}$. Thresholds $\tau_{c}=0.85$ and $\tau_{k}=1$ ensure strict adherence to cinematic principles.
Algorithm 1 Self-Validated Script Generation
1:User prompt $S$, shot count $N$, constraints $\mathcal{C}_{\text{film}}$
2:Initialize $\mathcal{P}\leftarrow\emptyset$, $p_{\text{prev}}\leftarrow\emptyset$
3:Generate draft $S^{\prime}\leftarrow\{s_{1},...,s_{N}\}=\mathcal{M}_{\text{LLM}}(S,N)$
4:for $i\leftarrow 1$ to $N$ do
5:     repeat
6:         $p_{i}^{\prime}\leftarrow\mathcal{M}_{\text{LLM}}(s_{i},\mathcal{C}_{\text{film}},p_{\text{prev}})$
7:         $C_{i}\leftarrow\mathbb{I}[C(p_{i},p_{i-1})>\tau_{c}]$
8:         $K_{i}\leftarrow\mathbb{I}[K(p_{i},\mathcal{C}_{\text{film}})>\tau_{k}]$
9:         $\mathcal{V}(p_{i})\leftarrow C_{i}\cdot K_{i}$
10:     until $\mathcal{V}(p_{i}^{\prime})=1$
11:     $\mathcal{P}\leftarrow\mathcal{P}\cup p_{i}^{\prime}$, $p_{\text{prev}}\leftarrow p_{i}^{\prime}$
12:end for
Our dynamic storyline modeling transforms user prompts into shot sequences through a two-stage process with self-validation mechanisms that enforce narrative coherence by rejecting candidates violating cinematic principles (Algorithm 1).
4.2 Identity-Aware Cross-Shot Propagation
We resolve visual inconsistency through cross-shot propagation machanism, which maintains critical attributes (e.g., hairstyle, facial structure) while permitting narrative-driven variations (e.g., expression, aging). Using scripts $\mathcal{P}$, we generate keyframes with consistent visual identities through a two-stage process:
$\mathbf{I}=\{\mathbf{I}_{i}\}_{i=1}^{N}=\mathcal{F}(\mathcal{P},\Psi)$ (8)
where $\mathcal{F}$ represents our identity-preserving generation pipeline and $\Psi$ denotes the parameters of the character schema. For each shot script $p_{i}\in\mathcal{P}$, we extract:
$e^{T}_{i}=E_{\text{GLM}}(p_{i})\in\mathbb{R}^{d},\quad\mathcal{C}_{\text{char}}=\mathcal{M}_{\text{LLM}}(\mathcal{P})=\{c_{j}\}_{j=1}^{M}$ (9)
where $E_{\text{GLM}}$ is the text encoder and $\mathcal{C}_{\text{char}}$ contains $M$ identity descriptors (e.g., Young Mary, Elderly Mary when describing given scripts Mary’s life ). Identity-Preserving Portrait (IPP) tokens are synthesized through:
$\text{IPP}_{j}=\mathcal{M}_{I}(c_{j})\in\mathbb{R}^{H\times W\times 3},\quad e^{I}_{j}=E_{\text{CLIP}}(\text{IPP}_{j})\in\mathbb{R}^{d_{v}}$ (10)
where $\mathcal{M}_{I}$ is a pre-trained text-to-image model and $E_{\text{CLIP}}$ denotes CLIP [38]’s vision encoder. We inject identity features into diffusion via cross-attention:
$Q$ $=W_{Q}z_{t}\in\mathbb{R}^{n\times d_{k}}$ (11)
$K$ $=\lambda[W_{K}e^{T}_{i};W^{\prime}_{K}e^{I}_{j}]\in\mathbb{R}^{2n\times d_{k}}$ (12)
$V$ $=\lambda[W_{V}e^{T}_{i};W^{\prime}_{V}e^{I}_{j}]\in\mathbb{R}^{2n\times d_{v}}$ (13)
$\text{Attn}(Q,K,V)$ $=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (14)
where $z_{t}\in\mathbb{R}^{n\times d}$ is the latent in step $t$, $\lambda$ balances the influence of text / identity, and [;][;][ ; ] denotes concatenation. Keyframe generation integrates both modalities:
$I_{i}=\mathcal{D}\left(z_{T}\big{|}e^{T}_{i},e^{I}_{j(i)}\right),\quad j(i)=\mathcal{M}_{\text{LLM}}(p_{i},\mathcal{C}_{\text{char}})$ (15)
By anchoring IPP tokens in narrative-derived character descriptors $\mathcal{C}_{\text{char}}$ and integrating them via attention mechanisms, we achieve robust identity fidelity across shots.
@@ 4.2 Identity-Aware Cross-Shot Propagation
4.3 Adjacent Latent Transition Mechanisms
We address transition artifacts through latent-space noise management across shot boundaries. Given keyframes $\{I_{i}\}_{i=1}^{N}$ and script embeddings $\{e^{T}_{i}\}_{i=1}^{N}$, we generate shot-wise latents:
$Z_{i}=\mathcal{M}_{V}(e^{T}_{i},e^{I}_{i},\epsilon_{i})\in\mathbb{R}^{f\times c\times h\times w}$ (16)
where $e^{T}_{i}=E_{\text{GLM}}(s_{i})$ uses simplified shot description $s_{i}$ rather than detailed script $p_{i}$, $e^{I}_{i}$ is the keyframe embedding, and $\epsilon_{i}\sim\mathcal{N}(0,\mathbf{I})$ is the initial noise. Inspired by FIFO [30], we implement boundary-aware noise reset for cross-shot transitions:
$\epsilon_{\text{boundary}}$ $\sim\mathcal{N}(0,\beta_{i}\mathbf{I}),\quad\beta_{i}=\gamma\cdot(1-\frac{i}{N})$ (17)
$Z_{\text{final}}$ $=\mathcal{R}(Z_{1},Z_{2},\ldots,Z_{N})$ (18)
where $\beta_{i}$ controls noise magnitude at shot boundaries with scaling factor $\gamma$, and $\mathcal{R}$ denotes our reset function:
$\mathcal{R}(Z_{1},\ldots,Z_{N})=\left[Z_{1}^{1:f},Z_{2}^{1:f},\ldots,Z_{N}^{1:f}\right]$ (19)
For each transition between shots $i$ and $i+1$, we reset the diffusion process with:
$z_{i+1}^{0}=\epsilon_{\text{boundary}}+\alpha\cdot z_{i}^{f}$ (20)
where $\alpha\in[0,1]$ controls the temporal continuity and $z_{i}^{f}$ is the final latent frame of the shot $i$. The complete video generation becomes:
$V=\mathcal{D}\left(\bigcup_{i=1}^{N}Z_{i}\right)\in\mathbb{R}^{T\times 3\times H\times W}$ (21)
This mechanism decodes $\{Z_{i}\}_{i=1}^{N}$ into a coherent video $V\in\mathbb{R}^{T\times 3\times H\times W}$ through boundary reset operations, preserving both narrative flow and visual continuity across $T=N\times f$ frames without requiring additional training.
Table 1: Quantitative comparison with state-of-the-art T2V baselines. We compare average CLIP scores, and the average FC and SC scores within and across shots between VGoT and baseline models. We use $\mathbb{bold}$ to highlight the highest and $\underline{underline}$ for the second high.
Model CLIP $\uparrow$ WS-FC $\uparrow$ CS-FC $\uparrow$ WS-SC $\uparrow$ CS-SC $\uparrow$
EasyAnimate [53] 0.2402 0.4705 0.0268 0.7969 0.2037
CogVideo [54] 0.2477 0.6099 0.0222 0.7424 0.2069
VideoCrafter1 [7] 0.2478 0.3706 0.0350 0.7623 0.1867
VideoCrafter2 [8] 0.2529 0.5569 0.0686 0.7981 0.1798
VGoT 0.2557 0.8138 0.2688 0.9717 0.4276
@@ 4.3 Adjacent Latent Transition Mechanisms
Figure 4: Visual comparison of VGoT and baselines
4.4 Multi-Shot Evaluation Protocol
Existing evaluation frameworks predominantly focus on single-shot quality, leaving multi-shot assessment underspecified. We introduce a hierarchical protocol that quantifies both intra-shot and inter-shot consistency:
$\Omega_{\text{multi-shot}}=\{\Omega_{\text{WS-FC}},\Omega_{\text{CS-FC}},\Omega_{\text{WS-SC}},\Omega_{\text{CS-SC}}\}$ (22)
Within-Shot Face Consistency (WS-FC) measures identity preservation within temporal sequences:
$\Omega_{\text{WS-FC}}(V_{i})=\frac{1}{f-1}\sum_{j=1}^{f-1}\cos\langle F_{j}^{i},F_{j+1}^{i}\rangle$ (23)
where $F_{j}^{i}$ represents facial features from frame $j$ of shot $i$ extracted using InsightFace [39, 20, 16, 3, 2, 11, 12, 19, 9, 10].
Cross-Shot Face Consistency (CS-FC) evaluates identity preservation across shots:
$\Omega_{\text{CS-FC}}(V)=\frac{1}{N-1}\sum_{i=1}^{N-1}\frac{1}{n}\sum_{j=1}^{n}\cos\langle F_{f-j+1}^{i},F_{j}^{i+1}\rangle$ (24)
where $n=8$ in our implementation, comparing the last $n$ frames of shot $i$ with the first $n$ frames of shot $i+1$.
Within-Shot Style Consistency (WS-SC) quantifies stylistic coherence through VGG-19 [43] features:
$\Omega_{\text{WS-SC}}(V_{i})=\frac{1}{f-1}\sum_{j=1}^{f-1}\cos\langle S_{j}^{i},S_{j+1}^{i}\rangle$ (25)
where $S_{j}^{i}$ represents flattened VGG-19 [43] features from frame $j$ of shot $i$.
Cross-Shot Style Consistency (CS-SC) assesses style preservation between adjacent shots:
$\Omega_{\text{CS-SC}}(V)=\frac{1}{N-1}\sum_{i=1}^{N-1}\frac{1}{n}\sum_{j=1}^{n}\cos\langle S_{f-j+1}^{i},S_{j}^{i+1}\rangle$ (26)
These metrics establish a comprehensive framework capturing both local and global consistency aspects essential for multi-shot video assessment, addressing limitations in current evaluation approaches.
@@ 4.4 Multi-Shot Evaluation Protocol
CLIP average $\uparrow$ PSNR $\uparrow$ IS $\uparrow$ WS-FC $\uparrow$ CS-FC $\uparrow$ WS-SC $\uparrow$ CS-SC $\uparrow$
w/o DSM w/o IPP 0.1146 24.3265 7.4624 0.7364 0.1129 0.9406 0.3650
w DSM w/o IPP 0.1146 24.3265 7.5783 0.7305 0.1174 0.9471 0.3663
w/o DSM w IPP 0.1223 23.9228 7.4521 0.8745 0.3291 0.9486 0.4186
Full Method 0.1111 25.7857 7.5194 0.8303 0.2738 0.9487 0.3859
@@ 5.1 Experiment Settings
Current video datasets lack sufficient multi-shot narratives with consistent characters across scenes. We therefore constructed a benchmark dataset using VGoT to create ten 30-shot stories (300 shots in total) for evaluation. Each story originates from a single user input $S$, which generates shot outlines $S^{\prime}$ and detailed scripts $\mathcal{P}$ across 30 shots and five domains defined in Eq. 6. For quantitative assessment, we employ four key metrics: Within-Shot Face Consistency ($\Omega_{\text{WS-FC}}$), Cross-Shot Face Consistency ($\Omega_{\text{CS-FC}}$), Within-Shot Style Consistency ($\Omega_{\text{WS-SC}}$), and Cross-Shot Style Consistency ($\Omega_{\text{CS-SC}}$). We additionally report CLIP score [38], PSNR [14], and IS [4]. Our implementation uses multiple GPT-4o [1] for scripting, Kolor [47] as base model for keyframes, and DynamiCrafter [52] for video generation, compared against EasyAnimate [53], CogVideo [27], and VideoCrafter [7, 8] on NVIDIA H100 GPUs.
5.2 Comparison Evaluation
Our evaluation compares VGoT against state-of-the-art text-to-video models using narrative scenarios from Sec 5.1. Quantitative results in Table 1 demonstrate VGoT’s superiority as follows.
VGoT achieves 0.8138 $\Omega_{\text{WS-FC}}$ and 0.9717 $\Omega_{\text{WS-SC}}$, outperforming the best baselines (VideoCrafter2’s 0.5569 $\Omega_{\text{WS-FC}}$ and 0.7981 $\Omega_{\text{WS-SC}}$) by 46.1% and 21.7% respectively. For cross-shot consistency, VGoT’s 0.2688 $\Omega_{\text{CS-FC}}$ and 0.4276 $\Omega_{\text{CS-SC}}$ surpass VideoCrafter1’s second-best 0.0686 $\Omega_{\text{CS-FC}}$ and CogVideo’s 0.2069 $\Omega_{\text{CS-SC}}$ by 291.8% and 106.7%. While outperform text-visual alignment via best CLIP score (0.2557), VGoT maintains this performance while requiring 30$\times$ less manual input.
Human evaluations (Table 2) confirm these findings: VGoT receives 66.7% ”Good” ratings for cross-shot consistency versus 27.2% for VideoCrafter2 and 23.2% for CogVideo. In visual quality, 70.0% of evaluators rate VGoT’s outputs as ”Good” compared to 52.4% for CogVideo. This preference is qualitatively validated in Fig 4, which demonstrates VGoT’s superior maintenance of character consistency and visual coherence across extended narratives compared to baseline outputs.
Figure 5: Visual Demonstration of the ablation studies of VGoT